### v1

@@ Abstract
It is well-established that an open continuum of electromagnetic modes can cause an excited atom to undergo radiative decay through spontaneous emission [1]. Careful engineering of the electromagnetic environment using cavities and waveguides can be used to control the spontaneous emission rate via the well-known Purcell effect [2, 3]. In cavity quantum electrodynamics (QED) and modern quantum science, this capability is harnessed at the single quantum level to control interactions between atoms and photons [4]. Representative examples include Josephson junction qubits coupled to microwave resonators in superconducting quantum processors [5], atoms and ions in Fabry-Perot cavities [6], and solid-state optical emitters in nanophotonic devices [7, 8, 9, 10] for quantum networking. The physics of an atom in an open continuum also applies to the emission of acoustic waves from an artificial atom in a solid, and is known to limit spin relaxation times in several color-centers [11, 12, 13, 14] and rare-earth ion impurities [15, 16] at low temperatures.
@@ 2 Results
Our devices consist of a rectangular cross-section diamond waveguide with a one-dimensional array of eye-shaped air holes (4.12). The unit cell dimensions and periodicity of the hole array are chosen to support a simultaneous photonic and phononic bandgap. These devices (Fig. 1f and Fig. 1g) are designed to support a high Q mechanical breathing mode and a high Q optical mode. Fig. 1d and Fig. 1e depict the simulated electric field magnitude for the TE optical mode (orange) and simulated strain and deformation profile (blue) for the device’s mechanical breathing mode, respectively. Light is coupled into the cavity region via a tapered fiber - diamond waveguide interface [43, 7] (Fig. 1f). The resonant frequency of the optical mode can be frequency-tuned into resonance with the SiV zero-phonon line (ZPL) at 737 nm (406 THz) resulting in an efficient spin-photon interface. This capability enables probing the spin state using ultra-low optical probe powers on the order of , thus limiting the laser-induced heating of the mechanical modes that has plagued previous experiments on SiV centers in optomechanical crystals [37]. Due to the significant mode overlap between the TE optical mode and the mechanical breathing mode of our device, there is a significant optomechanical coupling (4.6) due primarily to the photoelastic effect [39]. This allows for efficient probing of the properties of the mechanical mode () directly using optomechanical spectroscopy [36, 39, 37], and independently of spin-phonon coupling.
After the fabrication of the diamond optomechanical resonators, SiV centers are incorporated at the center of the optomechanical resonator by means of masked ion implantation [7, 44]. The implantation dose is chosen to result in 2-3 SiV centers on average per device. SiV centers in the same device are spectrally distinguishable from each other because of variations in the local strain environment of the diamond crystal lattice, which leads to variations in the optical transition frequencies. We address single SiV centers individually in optical frequency space by tuning a laser into resonance with the desired transition.
@@
To define a spin qubit with the SiV center, we apply a magnetic field of 2.5 kG along the z- axis in the lab frame ( with respect to the SiV center’s high symmetry axis) and lift the spin degeneracy. The frequencies of the Zeeman-split optical transitions, measured via resonant laser excitation spectroscopy, are used to estimate the spin transition frequency, for any given magnetic field condition (4.1). We initialize the SiV spin state by optically pumping on one of the Zeeman-split transitions. After initializing the spin into the state, the population decay rate is proportional to the acoustic density of states of the surrounding structure at . For the range of magnetic fields and the system temperature in these measurements, the spin population is expected to be predominantly in the state in the steady state, since . In the SI 4.8, we compare spin relaxation measurements performed with optical pumping into the and states and infer that the SiV environment is at a temperature of mK. At this temperature, the acoustic modes in the environment are close to their quantum ground state, and the GHz breathing mode is expected to have a thermal occupancy of 0.02 quanta.
@@
For the remainder of the discussion, we present data measured on the device using the ALD cavity-tuning method and minimal residual gas condensation, with corresponding data for the gas-tuning method shown in the SI 4.3. Fig. 3c shows the spin population decay curves measured at zero detuning from the breathing mode (solid curve) and with a detuning of -, clearly illustrating the Purcell enhancement in spin relaxation between the on and off-resonant cases. The optical pulse sequence used to measure the SiV center’s spin population decay rate at a given spin transition frequency is shown in the inset. A repump laser excites the f↑↑′ (f↓↓′) transition while a pump/probe laser tracks the f↓↓′ (f↑↑′). When the time delay between the pump and probe pulses is increased, the SiV center’s spin population read out using the probe decays towards its thermal equilibrium value. The data fits well to an exponential decay, allowing us to extract the spin relaxation rate. The Purcell-enhanced spin-decay rate due to a near ground state mechanical resonance where is given by where is the spin relaxation due to the environment of the spin mechanical resonator system. We measure kHz at a detuning of -0.1 GHz from the resonance which we take as the value of , and we measure kHz on resonance. With the linewidth of the resonance measured as MHz, we infer a spin-phonon coupling rate of kHz. Our simulations suggest that under ideal conditions of low static strain and accurate SiV placement at the breathing mode strain maximum, could attain a maximum value of 9 MHz (4.13). The difference between this maximum value and the experimentally measured value is well explained by the non-ideal conditions in the studied device, including the presence of static strain (4.2) and the typical expected error in SiV placement. Our measurements correspond to a -based cooperativity . Assuming the previously measured value for the SiV center spin linewidth, [7] in photonic crystal cavities with similar dimensions, this corresponds to a *-based cooperativity .
@@
In addition to the resonance induced by the mechanical breathing mode, several distinct features are observed in the spin decay spectrum of Fig. 4a. This broadband spectrum is taken with a aligned magnetic field. We attribute these resonances in the spin decay rate to other localized mechanical modes of the nanostructure with appreciable spin-phonon coupling to the SiV. Although these modes cannot be observed independently using optomechanical spectroscopy, numerical simulations of the acoustic mode structure of our device reveal several acoustic modes with appreciable spin-phonon coupling and negligible optomechanical coupling. This is because while the former simply requires an appreciable strain amplitude due to the mode at the SiV-center’s location, the latter requires an appreciable value for the photoelastic overlap integral between the optical and mechanical mode profiles extending over the volume of the structure. We find that our numerical simulations can reproduce the overall qualitative shape of the experimentally measured spin decay spectrum (4.14), confirming that our spin relaxation spectroscopy technique can detect phonon modes of the nanostructure beyond the resonances with appreciable optomechanical interactions (4.12). However, the precise frequencies of acoustic resonances detected (besides the breathing mode) and the corresponding magnitudes of Purcell enhancement are strongly dependent on the position of the SiV center within the OMC device (see supplementary information 4.13).
@@ 3 Conclusions and Outlook
In conclusion, we observed the acoustic Purcell effect between a single color-center spin qubit and a nanomechanical resonator. We demonstrated sensing of the broadband phonon spectrum of the nanostructure using the color-center as an atomic-scale acoustic probe. At the 12-GHz acoustic resonance of the device, we observe a Purcell enhancement or -based spin-phonon cooperativity, and a *- based cooperativity, . While this is the highest spin-phonon cooperativity observed to date, we expect future advances with our experimental platform to enable , the regime for most applications involving quantum-coherent interactions. In the present work, the cooperativity is primarily limited by excess mechanical damping (increased ) from the material deposition (gas condensation or ALD alumina) required to tune the co-localized optical mode into resonance with the SiV. An increase in cooperativity of at least two orders of magnitude is feasible with tuning methods that avoid material deposition, such as electrostatically-tunable zipper cavities [49]. In this regime, the system can be used to perform bi-directional transduction of quantum states between the spin qubit and the acoustic mode. The acoustic resonator in our platform is amenable to co-design with recently developed electro-optomechanical [20] and piezo-optomechanical quantum transducers [23, 50, 26], strengthening prospects for interconnecting color-center spin-based quantum memories with superconducting qubits. Such a hybrid quantum system would combine the exceptional coherence and atomic-scale footprint of color-center spin qubits with fast, high-fidelity superconducting qubit gates, and provide unique opportunities for next-generation quantum networks [51] and distributed quantum information processing [52].
@@
4 Methods
@@ 4.1 Calibration of Spin Transition Frequency
For our spin decay spectroscopy measurements, we use the and transitions because of their large photon scattering rate and comparatively small movement as a function of the applied magnetic field magnitude. The value of the conversion constant depends on the angle of the applied magnetic field and is calibrated at every angle for the angle-dependent measurements in Fig. 4. In particular, for a given field magnitude, a magnetic field oriented perpendicular to the SiV center’s symmetry axis will produce the minimum spin transition frequency, whereas a field aligned with the symmetry axis will produce the maximum spin transition frequency. The value of the conversion constant also depends on the static crystal strain experienced by the SiV center, which can be inferred from the ground state orbital splitting [11]. It was observed to change after ALD alumina cladding 4.2, and to a lesser extent after repeated gas tuning. Therefore, this calibration curve is retaken after any material deposition performed to tune the optical resonance.
@@ 4.3 Gas Tuning and ALD Cladding
The optical mode of our device can be tuned into resonance with the SiV center’s C-line by in-situ gas condensation. Gas is delivered to the sample through a copper tube thermally connected to the 4K plate of the dilution fridge. This tube inside the fridge is mated to a stainless steel tube which runs outside of the fridge. Resistive heaters and dedicated thermometers are installed on the gas tube allowing control of the tube temperature. During the cooldown, after the fridge reaches a temperature of about 4K, the tube is heated above the boiling temperature of inert gases, which can then condense on the sample. We observed that once the gas tube reaches a temperature in the range of 35-45K, the optical resonance of our device begins to red-shift, indicating gas condensation. Using this method, we achieve a coarse wavelength shift of the cavity resonance to a longer wavelength with respect to the SiV C line. After this step, a fine-tuning step is performed to blue-shift the cavity resonance onto the SiV C line. This is achieved via controlled boiling of the condensed gas with a high-power broadband laser (NKT Photonics, Super K supercontinuum white light laser) through the tapered fiber interface, which blue-shifts the resonance. During this tuning process, the optical resonance frequency is monitored in real time on a spectrometer (Princeton Instruments HRS-750).
@@ 4.5 Fabrication of Rectangular Cross-Section SiV-OMC Devices
The rectangular cross-section SiV-OMC devices are fabricated based on quasi-isotropic etching. [58, 37] An overview of the fabrication process is shown in Fig. 10. We use plasma-enhanced chemical vapor deposition to make a 100 nm-thick SiN layer on the bulk diamond before spin-coating a 400 nm-thick electron beam (EB) resist (ZEP 520A) [(Fig. 9(i))]. The OMC pattern is created on the EB resist by EB lithography and resist development [(Fig. 9(ii))]. The SiN layer is etched by inductively coupled plasma-reactive ion etching (ICP-RIE) [(Fig. 9(iii))]. After the EB resist is removed[(Fig. 9(iv))], the OMC pattern is transferred into the diamond by oxygen plasma RIE[(Fig. 9(v))]. We deposit 20 nm-thick Al2O3 by atomic layer deposition (ALD) for conformal coverage of the sample[(Fig. 9(vi))]. We etch the Al2O3 layer by RIE with Ar and Cl gases, maintaining the sidewalls of the OMC devices covered with Al2O3 [(Fig. 9(vii))]. We realize the free-standing structures using the quasi-isotropic etching technique based on an oxygen-based RIE[(Fig. 9(ix))]. We remove the Al2O3 and SiN layers by hydrofluoric acid.
To implant SiV centers in the fabricated OMC devices, we spin-coat a PMMA layer on the fabricated devices [(Fig. 9(xi))] and create the mask for the deterministic ion implantation process. After masked ion implantation [(Fig. 9(xii))], we remove the PMMA layer to complete the fabrication of OMC devices [(Fig. 9(xiii))]. The SiV centers in OMC devices are generated by annealing. The details of the mask ion implantation technique can be found in our previous report.[59] We note that our fabricated devices typically have a few SiV centers, we use an OMC device with a single SiV center in this work.

### v2

@@ Abstract
It is well-established that an open continuum of electromagnetic modes can cause an excited atom to undergo radiative decay through spontaneous emission [1]. Careful engineering of the electromagnetic environment using cavities and waveguides can be used to control the spontaneous emission rate via the well-known Purcell effect [2, 3]. In cavity quantum electrodynamics (QED) and modern quantum science, this capability is harnessed at the single quantum level to control interactions between atoms and photons [4]. Representative examples include Josephson junction qubits coupled to microwave resonators in superconducting quantum processors [5], atoms and ions in Fabry-Perot cavities [6], solid-state spins in microwave cavities [7], and solid-state optical emitters in nanophotonic devices [8, 9, 10, 11] for quantum networking. The physics of an atom in an open continuum also applies to the emission of acoustic waves from an artificial atom in a solid, and is known to limit spin relaxation times in several color-centers [12, 13, 14, 15] and rare-earth ion impurities [16, 17] at low temperatures.
@@ 2 Results
Our devices consist of a rectangular cross-section diamond waveguide with a one-dimensional array of eye-shaped air holes (5.12). The unit cell dimensions and periodicity of the hole array are chosen to support a simultaneous photonic and phononic bandgap. These devices (Fig. 1f and Fig. 1g) are designed to support a high Q mechanical breathing mode and a high Q optical mode. Fig. 1d and Fig. 1e depict the simulated electric field magnitude for the TE-like optical mode (orange) and simulated strain and deformation profile (blue) for the device’s mechanical breathing mode, respectively. Light is coupled into the cavity region via a tapered fiber - diamond waveguide interface [44, 8] (Fig. 1f). The resonant frequency of the optical mode can be frequency-tuned into resonance with the SiV zero-phonon line (ZPL) at 737 nm (406 THz) resulting in an efficient spin-photon interface. This capability enables probing the spin state using ultra-low optical probe powers on the order of , thus limiting the laser-induced heating of the mechanical modes that has plagued previous experiments on SiV centers in optomechanical crystals [38]. Due to the significant mode overlap between the TE-like optical mode and the mechanical breathing mode of our device, there is a significant optomechanical coupling (5.6) due primarily to the photoelastic effect [40]. This allows for efficient probing of the properties of the mechanical mode () directly using optomechanical spectroscopy [37, 40, 38], and independently of spin-phonon coupling.
After the fabrication of the diamond optomechanical resonators, which utilizes a quasi-isotropic diamond etch [45] SiV centers are incorporated at the center of the optomechanical resonator by means of masked ion implantation [8, 46]. The implantation dose is chosen to result in 2-3 SiV centers on average per device. SiV centers in the same device are spectrally distinguishable from each other because of variations in the local strain environment of the diamond crystal lattice, which leads to variations in the optical transition frequencies. We address single SiV centers individually in optical frequency space by tuning a laser into resonance with the desired transition.
@@
To define a spin qubit with the SiV center, we apply a magnetic field of 2.5 kG along the z- axis in the lab frame ( with respect to the SiV center’s high symmetry axis) and lift the spin degeneracy. The frequencies of the Zeeman-split optical transitions, measured via resonant laser excitation spectroscopy, are used to estimate the spin transition frequency, for any given magnetic field condition (5.1). We initialize the SiV spin state by optically pumping on one of the Zeeman-split transitions. After initializing the spin into the state, the population decay rate is proportional to the acoustic density of states of the surrounding structure at . For the range of magnetic fields and the system temperature in these measurements, the spin population is expected to be predominantly in the state in the steady state, since . In the SI 5.8, we compare spin relaxation measurements performed with optical pumping into the and states and infer that the SiV environment is at a temperature of mK. At this temperature, the acoustic modes in the environment are close to their quantum ground state, and the GHz breathing mode is expected to have a thermal occupancy of 0.02 quanta.
@@
For the remainder of the discussion, we present data measured on the device using the ALD cavity-tuning method and minimal residual gas condensation, with corresponding data for the gas-tuning method shown in the SI 5.3. Fig. 3c shows the spin population decay curves measured at zero detuning from the breathing mode (solid curve) and with a detuning of -, clearly illustrating the Purcell enhancement in spin relaxation between the on and off-resonant cases. The optical pulse sequence used to measure the SiV center’s spin population decay rate at a given spin transition frequency is shown in the inset. A repump laser excites the f↑↑′ (f↓↓′) transition while a pump/probe laser tracks the f↓↓′ (f↑↑′). When the time delay between the pump and probe pulses is increased, the SiV center’s spin population read out using the probe decays towards its thermal equilibrium value. The data fits well to an exponential decay, allowing us to extract the spin relaxation rate. The Purcell-enhanced spin-decay rate due to a near ground state mechanical resonance where is given by where is the spin relaxation due to the environment of the spin mechanical resonator system. We measure kHz at a detuning of -0.1 GHz from the resonance which we take as the value of , and we measure kHz on resonance. With the linewidth of the resonance measured as MHz, we infer a spin-phonon coupling rate of kHz. Our simulations suggest that under ideal conditions of low static strain and accurate SiV placement at the breathing mode strain maximum, could attain a maximum value of 9 MHz (5.13). The difference between this maximum value and the experimentally measured value is well explained by the non-ideal conditions in the studied device, including the presence of static strain (5.2) and the typical expected error in SiV placement. Our measurements correspond to a -based cooperativity . Assuming the previously measured value for the SiV center spin linewidth, [8] in photonic crystal cavities with similar dimensions, this corresponds to a *-based cooperativity .
@@
In addition to the resonance induced by the mechanical breathing mode, several distinct features are observed in the spin decay spectrum of Fig. 4a. This broadband spectrum is taken with a aligned magnetic field. We attribute these resonances in the spin decay rate to other localized mechanical modes of the nanostructure with appreciable spin-phonon coupling to the SiV. Although these modes cannot be observed independently using optomechanical spectroscopy, numerical simulations of the acoustic mode structure of our device reveal several acoustic modes with appreciable spin-phonon coupling and negligible optomechanical coupling. This is because while the former simply requires an appreciable strain amplitude due to the mode at the SiV-center’s location, the latter requires an appreciable value for the photoelastic overlap integral between the optical and mechanical mode profiles extending over the volume of the structure. We find that our numerical simulations can reproduce the overall qualitative shape of the experimentally measured spin decay spectrum (5.14), confirming that our spin relaxation spectroscopy technique can detect phonon modes of the nanostructure beyond the resonances with appreciable optomechanical interactions (5.12). However, the precise frequencies of acoustic resonances detected (besides the breathing mode) and the corresponding magnitudes of Purcell enhancement are strongly dependent on the position of the SiV center within the OMC device (see supplementary information 5.13).
@@ 3 Conclusions and Outlook
In conclusion, we observed the acoustic Purcell effect between a single color-center spin qubit and a nanomechanical resonator. We demonstrated sensing of the broadband phonon spectrum of the nanostructure using the color-center as an atomic-scale acoustic probe. At the 12-GHz acoustic resonance of the device, we observe a Purcell enhancement or -based spin-phonon cooperativity, and a *- based cooperativity, . While this is the highest spin-phonon cooperativity observed to date, we expect future advances in our experimental platform to enable , the regime for most applications involving quantum-coherent interactions. In the present work, the cooperativity is primarily limited by excess mechanical damping (increased ) from the material deposition (gas condensation or ALD alumina) required to tune the co-localized optical mode into resonance with the SiV. An increase in cooperativity of at least two orders of magnitude is feasible with tuning methods that avoid material deposition, such as electrostatically-tunable zipper cavities [51]. In this regime, the system can be used to perform bi-directional transduction of quantum states between the spin qubit and the acoustic mode. The acoustic resonator in our platform is amenable to co-design with recently developed electro-optomechanical [21] and piezo-optomechanical quantum transducers [24, 52, 27], strengthening prospects for interconnecting color-center spin-based quantum memories with superconducting qubits[53]. Such a hybrid quantum system would combine the exceptional coherence and atomic-scale footprint of color-center spin qubits with fast, high-fidelity superconducting qubit gates, and provide unique opportunities for next-generation quantum networks [54] and distributed quantum information processing [55].
@@
3.1 Data Availability
The data that support the findings in this study are available from the authors on reasonable request.
4 Acknowledgments
The authors would like to thank M. Bhaskar, D. Assumpcao, and C. Knaut for fruitful discussions. This research was supported by the National Science Foundation under grant number DMR-1231319, titled “Center for Integrated Quantum Systems (CIQM),” the Army Research Office/Department of the Army under award number W911NF1810432, titles “Ab-Initio Solid-State Quantum Materials: Design, Production, and Characterization at the Atomic Scale,” the Office of Naval Research under award number N00014-20-1-2425, titled “Quantum Information Processing With Phonons,” the National Science Foundation under award number EEC-1941583, titled “NSF Engineering Research Center for Quantum Networks (CQN),” the Air Force Office of Scientific Research under award number FA9550-23-1-0333, titled “Quantum Phononics to Advance Quantum Information Processing,” the National Research Foundation funded by the Korean government (NRF-2022M3K4A1094782), and Amazon Web Services under award number A50791, titled “Partnership for Quantum Networking.” G. J was supported in part by the Natural Sciences and Research Council of Canada (NSERC). K.K. acknowledges financial support from JSPS Overseas Research Fellowships (Project No. 202160592). This work was partly supported by the KIST institutional program (2E33541, 2E33571) funded by the Korea Institute of Science and Technology, and the National Research Foundation of Korea (NRF) grant (No. RS-2022-NR068818), funded by the Korean government (MSIT). C. C was supported in part by Singapore’s Agency for Science, Technology and Research (A*STAR). H. W acknowledges financial support from the NSF GRFP fellowship. B. P. acknowledges financial support from Q-NEXT, a U.S. Department of Energy Office of Science National Quantum Information Science Research Centers under award number DE-FOA-000225. B.M. is involved in developing diamond photonic technology at Lightsynq Technologies Inc. This work was performed in part at the Center for Nanoscale Systems (CNS), a member of the National Nanotechnology Infrastructure Network (NNIN), which is supported by the National Science Foundation award ECS-0335765. CNS is part of Harvard University.
5 Supplementary information
@@ 5.1 Calibration of Spin Transition Frequency
For our spin decay spectroscopy measurements, we use the and transitions because of their large photon scattering rate and comparatively small movement as a function of the applied magnetic field magnitude. The value of the conversion constant depends on the angle of the applied magnetic field and is calibrated at every angle for the angle-dependent measurements in Fig. 4. In particular, for a given field magnitude, a magnetic field oriented perpendicular to the SiV center’s symmetry axis will produce the minimum spin transition frequency, whereas a field aligned with the symmetry axis will produce the maximum spin transition frequency. The value of the conversion constant also depends on the static crystal strain experienced by the SiV center, which can be inferred from the ground state orbital splitting [12]. It was observed to change after ALD alumina cladding 5.2, and to a lesser extent after repeated gas tuning. Therefore, this calibration curve is retaken after any material deposition performed to tune the optical resonance.
@@ 5.3 Gas Tuning and ALD Cladding
The optical mode of our device can be tuned into resonance with the SiV center’s C-line by in-situ gas condensation. Gas is delivered to the sample through a copper tube thermally connected to the 4K plate of the dilution refrigerator. This tube inside the fridge is mated to a stainless steel tube which runs outside of the fridge. Resistive heaters and dedicated thermometers are installed on the gas tube, allowing control of the tube temperature. During the cooldown, after the fridge reaches a temperature of about 4K, the tube is heated above the boiling temperature of inert gases, which can then condense on the sample. We observed that once the gas tube reaches a temperature in the range of 35-45K, the optical resonance of our device begins to red-shift, indicating gas condensation. Using this method, we achieve a coarse wavelength shift of the cavity resonance to a longer wavelength with respect to the SiV C line. After this step, a fine-tuning step is performed to blue-shift the cavity resonance onto the SiV C line. This is achieved via controlled boiling of the condensed gas with a high-power broadband laser (NKT Photonics, Super K supercontinuum white light laser) through the tapered fiber interface, which blue-shifts the resonance. During this tuning process, the optical resonance frequency is monitored in real time on a spectrometer (Princeton Instruments HRS-750).
@@ 5.5 Fabrication of Rectangular Cross-Section SiV-OMC Devices
The rectangular cross-section SiV-OMC devices are fabricated by quasi-isotropic etching. [45, 61, 38] An overview of the fabrication process is shown in Fig. 10. We use plasma-enhanced chemical vapor deposition to make a 100 nm-thick SiN layer on the bulk diamond before spin-coating a 400 nm-thick electron beam (EB) resist (ZEP 520A) [(Fig. 9(i))]. The OMC pattern is created on the EB resist by EB lithography and resist development [(Fig. 9(ii))]. The SiN layer is etched by inductively coupled plasma-reactive ion etching (ICP-RIE) [(Fig. 9(iii))]. After the EB resist is removed[(Fig. 9(iv))], the OMC pattern is transferred into the diamond by oxygen plasma RIE[(Fig. 9(v))]. We deposit 20 nm-thick Al2O3 by atomic layer deposition (ALD) for conformal coverage of the sample[(Fig. 9(vi))]. We etch the Al2O3 layer by RIE with Ar and Cl gases, maintaining the sidewalls of the OMC devices covered with Al2O3 [(Fig. 9(vii))]. We realize the free-standing structures using the quasi-isotropic etching technique based on an oxygen-based RIE[(Fig. 9(ix))]. We remove the Al2O3 and SiN layers by hydrofluoric acid.
To implant SiV centers in the fabricated OMC devices, we spin-coat a PMMA layer on the fabricated devices [(Fig. 9(xi))] and create the mask for the deterministic ion implantation process. After masked ion implantation [(Fig. 9(xii))], we remove the PMMA layer to complete the fabrication of OMC devices [(Fig. 9(xiii))]. The SiV centers in OMC devices are generated by annealing. The details of the mask ion implantation technique can be found in our previous report[62]. We note that our fabricated devices typically have a few SiV centers, we use an OMC device with a single SiV center in this work.